OpenStudy (anonymous):

Find the indicated derivative using implicit differentiation (8xe^y)-(ye^x)=29 ; dy/dx

OpenStudy (bahrom7893):

(8x)e^y(dy/dx)+8e^y - (dy/dx)e^x-ye^x = 0 Solve for dy/dx

OpenStudy (bahrom7893):

(8x)e^y(dy/dx) - (dy/dx)e^x-ye^x = -8e^y

OpenStudy (bahrom7893):

(8x)e^y(dy/dx) - (dy/dx)e^x= -8e^y + ye^x (dy/dx)(8xe^y - e^x) = -8e^y + ye^x dy/dx = (-8e^y + ye^x )/(8xe^y - e^x)

OpenStudy (bahrom7893):

check the math

OpenStudy (anonymous):

take derivate and solve for y prime

OpenStudy (anonymous):

i can start writing it down,headace about to go home and sleep

OpenStudy (anonymous):

Yes! Thank you bahrom. I guess i have to study this some more. so confusing...

OpenStudy (anonymous):

Am getting a different answer here though...

OpenStudy (anonymous):

8e^y-e^xy

OpenStudy (anonymous):

move the constant over and took the derivate to be 0